Calculate the volume of 8.46×10-2 M potassium hydroxide required to neutralize 15.3 mL of a 0.172 M nitric acid solution.

________ mL
1
Expert's answer
2019-03-11T08:18:44-0400

V(KOH)*N(KOH) = V(HNO3)*N(HNO3)

V(KOH) = V(HNO3)*N(HNO3)/N(KOH)

V(KOH) = 15.3*0.172/0.0842 = 31.106 (ml(KOH))